# Kac–Moody Symmetric Spaces: arbitary symmetrizable complex or almost split real type

### v3

@@ 9 Symmetric Spaces of Almost Split Kac–Moody Groups
In this section, we going to define a symmetric space for almost split Kac–Moody $\mathbb{R}$-groups.
Therefore, let $G$ be a Kac–Moody group over $\mathbb{C}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix of size $n$ with rank $l$. Denote by $\Gamma$ the group of Galois automorphism induced by $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ on $G$ and assume that $G$ possess an algebraic $\mathbb{R}$-form $(\mathcal{F},\mathcal{U}_{\mathbb{R}})$ satisfying the (8.2) condition, i.e.
@@
By Section 8.1 1 one knows how $\overline{\sigma}$ acts on an element of a root group $U_{\alpha}$, $\alpha\in\Delta^{\mathrm{re}}$,:
@@
where we denote for simplicity reasons with $\Gamma$ the group of rectified and conjugated Galois automorphism $\widehat{\sigma}$.
It follows that $\Theta$ is a continuous involution on the topological group $G(\mathbb{R})$ and we define
@@
Since $\mathcal{F}(\mathbb{C})=G$ if $G$ is a complex (split) Kac–Moody group, it follows that the theorem is in this particular case just 4.4 for $\mathbb{K}=\mathbb{C}$. Furthermore, if we consider a real split Kac–Moody group in the theorem, the following theorem corresponds to 4.4 for $\mathbb{K}=\mathbb{R}$.
@@ 9.1 Flats in Almost Split Kac–Moody Symmetric Spaces
After establishing the notion of a symmetric space for almost split Kac–Moody groups, we can define flats. As we in the proof of 4.8, we define a symmetric Euclidean space in the fixed point set of the Cartan subalgebra and use the exponential map to obtain an isomorphism.
@@
be an element of the Cartan algebra, where $r_{i}\in\mathbb{C}$ are scalars and $v_{i}\in\Lambda^{\vee}$ are the $\mathbb{Z}$-basis of $\Lambda^{\vee}$ (cp. Section 3.2). Now we can compute
@@
Recall from 3.20 that an element $g\in AK$ can be written as $ak$ for a unique $a\in A$ and a unique $k\in K$. Assume that $ak\in AK\cap G(\mathbb{R})\subseteq G$ and let $\overline{\sigma}$ a rectified Galois automorphism. Then it follows that $\overline{\sigma}(A)\subseteq A$ and $\overline{\sigma}(K)\subseteq K$. From the condition (8.2) it further follows that
@@
In the proof of the characterization, one uses that the group model of the symmetric spaces is isomorphic to the coset model. But, the isomorphism between the coset model and the group model established in Section 4.3 is independent of the Kac–Moody group. It is essentially based on the isomorphism specified in 2.4 for abstract symmetric spaces according to Loos and on the existence of the twist map $\tau$. Hence, it follows that
@@
From [AB08, Proposition 6.27] it follows that the stabilizer of a spherical residue is equal to a spherical parabolic subgroup. From [Rém02, Definition 12.1.3] follows that a parabolic subgroup in $G(\mathbb{R})$ is equal to a parabolic subgroup in $G$. Using the same arguments as in the last step of the proof of 4.1, it follows that this can be interpreted as a subgroup of $\operatorname{GL}_{m}(\mathbb{R})$, $m\in\mathbb{N}$.

### v4

@@ 9 Symmetric Spaces of Almost Split Kac–Moody Groups
In this section, we are going to define a symmetric space for almost split Kac–Moody $\mathbb{R}$-groups.
Therefore, let $G$ be a Kac–Moody group over $\mathbb{C}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix of size $n$ with rank $l$. Denote by $\Gamma$ the group of Galois automorphisms induced by $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ on $G$ and assume that $G$ possesses an algebraic $\mathbb{R}$-form $(\mathcal{F},\mathcal{U}_{\mathbb{R}})$ satisfying the (8.2) condition, i.e.
@@
By Section 8.1 1 one knows how $\overline{\sigma}$ acts on an element of a root group $U_{\alpha}$, $\alpha\in\Delta^{\mathrm{re}}$:
@@
where we denote with $\Gamma$ the group of rectified and conjugated Galois automorphism $\widehat{\sigma}$. Note that this shows that $G(\mathbb{R})$ is closed under the action of $\Theta$.
Therefore $\Theta$ is a continuous involution on the topological group $G(\mathbb{R})$ and we define
@@
Since $\mathcal{F}(\mathbb{C})=G$ if $G$ is a complex (split) Kac–Moody group, it can be deduced that the theorem is in this particular case just 4.4 for $\mathbb{K}=\mathbb{C}$. Furthermore, if we consider a real split Kac–Moody group in the theorem, the following theorem corresponds to 4.4 for $\mathbb{K}=\mathbb{R}$.
@@ 9.1 Flats in Almost Split Kac–Moody Symmetric Spaces
After establishing the notion of a symmetric space for almost split Kac–Moody groups, we can define flats. As we have seen in the proof of 4.8, we define a symmetric Euclidean space in the fixed point set of the Cartan subalgebra and use the exponential map to obtain an isomorphism.
@@
be an element of the Cartan algebra, where $r_{i}\in\mathbb{C}$ are scalars and $v_{i}\in\Lambda^{\vee}$ are the $\mathbb{Z}$-basis of $\Lambda^{\vee}$ (cf. Section 3.2). Now we can compute
@@
Recall from 3.20 that an element $g\in AK$ can be written as $ak$ for a unique $a\in A$ and a unique $k\in K$. Assume that $ak\in AK\cap G(\mathbb{R})\subseteq G$ and let $\overline{\sigma}$ a rectified Galois automorphism. Then we have $\overline{\sigma}(A)\subseteq A$ and $\overline{\sigma}(K)\subseteq K$. From the condition (8.2) it further follows that
@@
In the proof of the characterization, one uses that the group model of the symmetric spaces is isomorphic to the coset model. But, the isomorphism between the coset model and the group model established in Section 4.3 is independent of the Kac–Moody group. It is essentially based on the isomorphism specified in 2.4 for abstract symmetric spaces according to Loos and on the existence of the twist map $\tau$. Hence, as a result we have
@@
From [AB08, Proposition 6.27] follows that the stabilizer of a spherical residue is equal to a spherical parabolic subgroup. From [Rém02, Definition 12.1.3] follows that a parabolic subgroup in $G(\mathbb{R})$ is equal to a parabolic subgroup in $G$. Using the same arguments as in the last step of the proof of 4.1, it can be be interpreted as a subgroup of $\operatorname{GL}_{m}(\mathbb{R})$, $m\in\mathbb{N}$.